please show me how to do solve by addition I get confused when one equation is neg and the other is pos. like this problem
x+y=2
x-y=4

Answer by praseena(37)

x + y = 2
x - y = 4
Now add these two equations

x + y + x - y = 2 + 4
combune the like terms together.
x + x + y - y = 6
==> 2x = 6
now divide both sides of the equation by 2.
==> 2x/2 = 6/2
==> x = 3
thus we have got the value x = 3


now substitute the value of x in any of the given equation.
if we substitite it in the first equation,
we have, 3 + y = 2
==> y = 2-3
= -1

so solution of the equation is

x = 3 and
y = -1


Bye!!!!!!!!!!